### v3

@@ 1.1 Related work
Recently, Grapiglia and Nesterov [22] extended the accelerated CR method [37] for unconstrained smooth convex optimization to problem (1) with a twice differentiable convex $f$ whose Hessian is Hölder continuous, and achieved the global complexity bound of the objective values. For this class of problem (1), under an additional uniform convexity111 A function $h\!:\mathbb{X}\to\overline{\mathbb{R}}$ is said to be uniformly convex of degree $\alpha\geq 2$ over a convex set $C\subset{\rm dom}\,h$ if for some constant $\sigma>0$, $h(y)\geq h(x)+\langle\xi,y-x\rangle+\frac{\sigma}{\alpha}\|y-x\|^{\alpha}$ holds for all $x,y\in C$ and $\xi\in\partial h(x)$. of $f$ with degree $\nu\!+\!2$, where $\nu$ is the order of Hölder continuity, Doikov and Nesterov [18] established the linear convergence rate of objective value sequence of the CR method. Doikov and Nesterov [17, 20] also investigated (inexact) tensor methods for problem (1) with convex $f$ by the $p$th-order derivatives of $f$ with $p\geq 2$, which reduce to an (inexact) CR method for $p=2$. They proposed in [17] two dynamic strategies for choosing the computation accuracy of subproblems in terms of the number of iterations and the objective value difference of successive iterates, respectively, and for the latter scheme, achieved the global complexity of the objective values, and the linear convergence rate of objective value sequence under the uniform convexity of $F$ with degree $p+1$. When the inner accuracy becomes better, they also established superlinear convergence rate of order $(p\!+\!1)/2$ for the objective value sequence under the strong convexity of $F$. In [20], they proposed a regularized composite tensor method, and obtained the local $Q$-superlinear convergence of order $p/(\alpha\!-\!1)$ of the objective value sequence for $p>\alpha\!-\!1$ under the uniform convexity of $F$ with degree $\alpha\geq 2$ and the exact computation of subproblems. It is worth mentioning that Doikov and Nesterov [19] also considered a fully composite convex problem, covering problem (1) with convex $f$, and proved the linear convergence rate of objective value sequence under the uniform convexity of all component functions with degree $p+1$. When the uniform convexity is absent, they analyzed the global convergence of objective value sequence by requiring the outer function to be subhomogeneous.
@@
For nonconvex problem (1) where $f$ is $p$-times continuously differentiable with the $p$th-order derivative being Hölder continuous of degree $\beta_{p}\in[0,1]$, Cartis et al. [14] studied the worst-case complexity bound in terms of the first-order optimality conditions, and achieved the complexity bound $O(\epsilon^{-\frac{2+\beta_{2}}{1+\beta_{2}}})$ for the CR method. Necoara and Lupu [36] proposed a general higher-order majorization-minimization algorithm framework for minimizing a nonconvex and nonsmooth function by its $p$ higher-order surrogate at each iterate, and proved the superlinear convergence rate of the objective value sequence under the KL property with exponent $\theta\in(0,\frac{1}{1+p})$. Later, Nabou and Necoara [35] developed a general composite higher-order algorithmic framework for (1) with $f=\varphi\circ G$ and nonconvex $g$ by a $p$ higher-order surrogate of $G$ at each iterate. When higher-order composite surrogates are sufficiently smooth, they derived the global sublinear convergence rate of order $O(k^{-\frac{p}{p+1}})$ in terms of the first-order optimality conditions, and the $R$-linear convergence rate of the objective value sequence under the KL property of $F$ with exponent $\theta\in(0,\frac{p}{1+p}]$. The CR method for problem (1) precisely corresponds to the frameworks in [36, 35] with $p=2$.
@@
Meanwhile, the uniform convexity of degree $\alpha\geq 2$ or strong convexity of objective functions is very restrictive and implies that the associated optimization problems have a unique minimum. The local strong convexity of $F$ around a critical point is also restrictive, which implies that this critical point is an isolated local minimum of (1). Such a property is absent in many applications of interest; for example, the nonconvex quadratic minimization problem over a box constraint with the objective function of the form $F(x)=\frac{1}{2}x^{\top}Qx+c^{\top}x+\chi_{[l,u]}(x)$, where $Q$ is an $n\times n$ real symmetric matrix $Q$ with $\lambda_{\rm min}(Q)<0$ and $\chi_{[l,u]}(\cdot)$ is the indicator function over the box set $[l,u]\subset\mathbb{R}^{n}$. This nonconvex and nonsmooth problem does not possess local strong convexity around those stationary points that are not local minimum. In addition, the KL property of $F$ with exponent $\theta\in(0,{1}/{3})$ is also stringent because it is almost impossible to find such $F$ for problem (1) though there are examples for $F$ to be a KL function of exponent $1/2$ by [29, 51]. Recall that Yue et al. [49] established the local quadratic convergence rate of the iterate sequence of the CR method for unconstrained smooth optimization under a local Lipschitzian error bound. Then, it is natural to ask whether a practical inexact CR method can be designed for (1) to possess a quadratic convergence rate without the local strong convexity of $f$. This work aims to provide an affirmative answer to this question.
@@ 1.2 Main contribution
(1) For the proposed algorithm, we establish the full convergence of the iterate sequence under the KL property of $F$, by combining the sufficient decrease of the objective value sequence with the KL inequality on a potential function in a skillful way. Based on the full convergence analysis and the proof of [40, Theorem 3.2], when $F$ is a KL function of exponent $\theta\in\!(0,\frac{q-1}{q})$, we obtain the superlinear convergence rate of order $\frac{q-1}{\theta q}$ for the iterate and objective value sequences. These results are consistent with those of [40, Theorem 4.1 (ii)] for the algorithm of only theoretical values. From [3, Section 4] and [46], KL functions are ubiquitous in nonconvex and nonsmooth optimization; while KL property of exponent $\theta\in(\frac{1}{2},\frac{q-1}{q})$ for $q\in(2,3]$ is weaker than that of exponent $1/2$. Thus, our superlinear convergence results will have extensive application. Moreover, the superlinear convergence result on the objective value sequence also improves the one obtained in [36] with $p=2$ for problem (1), which is a special case of the optimization problem considered in [36].
@@
(2) Under a local Hölderian error bound of order $\gamma\in\!(\frac{1}{q-1},1]$ on a second-order stationary point set that reduces to the ordinary one used in [23, 49] if $g\equiv 0$, we prove that the iterate sequence converges to a second-order stationary point with a local $Q$-superlinear convergence rate of order $\gamma(q\!-\!1)$, and the convergence of the objective value sequence has the same superlinear rate under an additional lower second-order growth on $F$ around the second-order stationary point. This implies that the iterate and objective value sequences of our inexact CR method converge quadratically to a second-order stationary point and a second-order stationary value, respectively, under a local Lipschitzian error bound. This error bound condition is much weaker than the strong convexity assumption required in [20], and is basically the same as the one used in [49] when $g\equiv 0$. For the aforementioned quadratic minimization problem over a box set, though it does not possess the local strong convexity around any non-minimum stationary points, the local Lipschitzian error bound on the set of stationary points and the KL property of $F$ with exponent $1/2$ are implied by the local upper Lipschitz continuity of $(\partial F)^{-1}$ due to [41, Proposition 1]. To the best of our knowledge, this is the first practical inexact CR method for problem (1) to have a quadratic convergence rate without the local strong convexity of $F$, improving the superlinear convergence rate of order $4/3$ obtained in [40] under the KL property of $F$ with exponent $1/2$, and extending the work [49] to nonconvex and nonsmooth composite optimization. It is worth emphasizing that our proof of the local quadratic convergence rate for $q=3$ is not a direct extension of the proofs in [40, 49]. We also derive a global complexity bound $O(\epsilon^{-\frac{q}{q-1}})$ in terms of the first-order optimality conditions under the strict continuity of $\nabla^{2}\!f$ on a compact convex set containing the iterate sequences. When $g\equiv 0$ and $q=3$, this bound matches that of the CR method [38] and its variants [13] for unconstrained smooth optimization; when $g=\chi_{C}$ for a closed convex set $C\subset\mathbb{X}$, it agrees with the complexity bound in [14, Theorem 3.12].
@@
(3) We apply the proposed algorithm armed with ZeroFPR [44] as the solver of subproblems to three classes of problem (1), and compare its performance with that of ZeroFPR applied directly to these problems and IRPNM proposed in [31]. Numerical results show that our algorithm indeed has the performance matching theoretical results, and for the convex $\ell_{1}$-norm regularized logisitic regression, our CR method yields the optimal values and KKT residuals with better accuracy within $54$ iterations except for data “arcene”. For the nonconvex $\ell_{1}$-norm regularized Student’s t-regression, it returns the same objective values and the comparable KKT residuals within less iterations, and the returned solutions are checked to be second-order stationary points; and for the portfolio decision with higher moments, it yields the comparable objective values and better KKT residuals within much less running time than ZeroFPR and comparable running time with IRPNM.
@@ 2 Preliminaries
We first recall the strict continuity and subdifferential continuity of a proper function $h\!:\mathbb{X}\to\overline{\mathbb{R}}$. By [42, Definition 9.1], $h$ is strictly continuous at $\overline{x}\in{\rm dom}\,h$ if there exist a neighborhood $V$ of $\overline{x}$ and a constant $\kappa\geq 0$ such that for any $x^{\prime},x^{\prime\prime}\in V$, $|h(x^{\prime})-h(x^{\prime\prime})|\leq\kappa\|x^{\prime}-x^{\prime\prime}\|$; and it is strictly continuous at $\overline{x}$ relative to ${\rm dom}\,h$ if there exists a neighborhood $V$ of $\overline{x}$ such that $h$ is Lipschitz continuous on ${\rm dom}\,h\cap V$. By [42, Definition 13.28], the function $h$ is subdifferentially continuous at $\overline{x}$ for $\overline{v}$ if $\overline{v}\in\partial h(\overline{x})$ and, whenever $(x^{k},v^{k})\xrightarrow{{\rm gph}\,\partial h}(\overline{x},\overline{v})$, one has $h(x^{k})\to h(\overline{x})$; and when this holds for all $\overline{v}\in\partial h(\overline{x})$, it is said to be subdifferentially continuous at $\overline{x}$.
@@ 2.1 Stationary points of problem (1)
and denote by $\mathcal{X}^{*}$ the set of second-order stationary points. When $g\equiv 0$, this reduces to the one used in [49]. When $g\not\equiv 0$, for a given $x\in\mathcal{S}^{*}$, the following inclusion
$\mathcal{X}^{*}\supset\big{\{}x\in\mathcal{S}^{*}\,|\,\nabla^{2}\!f(x)\succeq 0\big{\}}$
for a nonconvex $f$ is generally strict because the set ${\rm prox}_{g}(\mathbb{X})-x$ is much smaller than $\mathbb{X}$. For example, when $g=\chi_{C}$ for a closed convex set $C\subset\mathbb{X}$, ${\rm prox}_{g}(\mathbb{X})-x\subset C-x$. When $g\not\equiv 0$ but $f$ is convex, $\mathcal{X}^{*}=\big{\{}x\in\mathcal{S}^{*}\,|\,\nabla^{2}\!f(x)\succeq 0\big{\}}=\mathcal{S}^{*}$.
@@ 2.2 Metric subregularity of residual mapping
(see [28, Definition 3.1]) Let $\mathcal{F}\!:\mathbb{X}\rightrightarrows\mathbb{X}$ be a multifunction. Consider any point $(\overline{x},\overline{y})\in{\rm gph}\,\mathcal{F}$, the graph of $\mathcal{F}$. For a given $\gamma>0$, we say that $\mathcal{F}$ is (metrically) $\gamma$-subregular at $(\overline{x},\overline{y})$ if there exist $\kappa>0$ and $\delta>0$ such that

### v4

@@ 1.1 Related work
Recently, Grapiglia and Nesterov [22] extended the accelerated CR method [36] for unconstrained smooth convex optimization to problem (1) with a twice differentiable convex $f$ whose Hessian is Hölder continuous, and achieved the global complexity bound of the objective values. For this class of problem (1), under an additional uniform convexity  [^1] A function $h\!:\mathbb{X}\to\overline{\mathbb{R}}$ is said to be uniformly convex of degree $\alpha\geq 2$ over a convex set $C\subset{\rm dom}\,h$ if for some constant $\sigma>0$, $h(y)\geq h(x)+\langle\xi,y-x\rangle+\frac{\sigma}{\alpha}\|y-x\|^{\alpha}$ holds for all $x,y\in C$ and $\xi\in\partial h(x)$. of $f$ with degree $\nu\!+\!2$, where $\nu$ is the order of Hölder continuity, Doikov and Nesterov [18] established the linear convergence rate of objective value sequence of the CR method. Doikov and Nesterov [17, 20] also investigated (inexact) tensor methods for problem (1) with convex $f$ by the $p$th-order derivatives of $f$ with $p\geq 2$, which reduce to an (inexact) CR method for $p=2$. They proposed in [17] two dynamic strategies for choosing the computation accuracy of subproblems in terms of the number of iterations and the objective value difference of successive iterates, respectively, and for the latter scheme, achieved the global complexity of the objective values, and the linear convergence rate of objective value sequence under the uniform convexity of $F$ with degree $p+1$. When the inner accuracy becomes better, they also established superlinear convergence rate of order $(p\!+\!1)/2$ for the objective value sequence under the strong convexity of $F$. In [20], they proposed a regularized composite tensor method, and obtained the local $Q$-superlinear convergence of order $p/(\alpha\!-\!1)$ of the objective value sequence for $p>\alpha\!-\!1$ under the uniform convexity of $F$ with degree $\alpha\geq 2$ and the exact computation of subproblems. It is worth mentioning that Doikov and Nesterov [19] also considered a fully composite convex problem, covering problem (1) with convex $f$, and proved the linear convergence rate of objective value sequence under the uniform convexity of all component functions with degree $p+1$. When the uniform convexity is absent, they analyzed the global convergence of objective value sequence by requiring the outer function to be subhomogeneous.
@@
For nonconvex problem (1) where $f$ is $p$-times continuously differentiable with the $p$th-order derivative being Hölder continuous of degree $\beta_{p}\in[0,1]$, Cartis et al. [14] studied the worst-case complexity bound in terms of the first-order optimality conditions, and achieved the complexity bound $O(\epsilon^{-\frac{2+\beta_{2}}{1+\beta_{2}}})$ for the CR method. Necoara and Lupu [35] proposed a general higher-order majorization-minimization algorithm framework for minimizing a nonconvex and nonsmooth function by its $p$ higher-order surrogate at each iterate, and proved the superlinear convergence rate of the objective value sequence under the KL property with exponent $\theta\in(0,\frac{1}{1+p})$. Later, Nabou and Necoara [34] developed a general composite higher-order algorithmic framework for (1) with $f=\varphi\circ G$ and nonconvex $g$ by a $p$ higher-order surrogate of $G$ at each iterate. When higher-order composite surrogates are sufficiently smooth, they derived the $R$-linear convergence rate of the objective value sequence under the KL property of $F$ with exponent $\theta\in(0,\frac{p}{1+p}]$. The CR method for problem (1) precisely corresponds to the frameworks in [35, 34] with $p=2$.
@@
Meanwhile, the uniform convexity of degree $\alpha\geq 2$ or strong convexity of objective functions is very restrictive and implies that the associated optimization problems have a unique minimum. The local strong convexity of $F$ around a critical point is also restrictive, which implies that the critical point is an isolated local minimum of (1). Such a property is absent in many applications of interest; for example, the nonconvex quadratic minimization problem over a box constraint with the objective function of the form $F(x)=\frac{1}{2}x^{\top}Qx+c^{\top}x+\chi_{[l,u]}(x)$, where $Q$ is an $n\times n$ real symmetric matrix $Q$ with $\lambda_{\rm min}(Q)<0$ and $\chi_{[l,u]}(\cdot)$ is the indicator function over the box set $[l,u]\subset\mathbb{R}^{n}$. This nonconvex and nonsmooth problem does not possess local strong convexity around those non-minimum stationary points. The KL property of $F$ with exponent $\theta\in(0,{1}/{3})$ is also stringent because it is almost impossible to find such $F$ for (1) though there are KL examples of exponent $1/2$ for $F$ by [28, 49]. Recall that Yue et al. [47] established the local quadratic convergence rate of the iterate sequence of the CR method for unconstrained smooth optimization under a local Lipschitzian error bound. It is natural to ask whether a practical inexact CR method can be designed for (1) to have a quadratic convergence rate under a similar local error bound. This work aims to provide an affirmative answer to the question.
@@ 1.2 Main contribution
(1) For the proposed algorithm, we establish the full convergence of the iterate sequence under the KL property of $F$ by combining the sufficient decrease of the objective values with the KL inequality on a potential function skillfully. Based on the full convergence analysis and the proof of [39, Theorem 3.2], under the KL property of $F$ with exponent $\theta\in\!(0,\frac{q-1}{q})$, we obtain the superlinear convergence rate of order $\frac{q-1}{\theta q}$ for the iterate and objective value sequences, which are consistent with those of [39, Theorem 4.1 (ii)] for the algorithm of only theoretical values. KL functions are ubiquitous in nonconvex and nonsmooth optimization by [3, Section 4], and KL property of exponent $\theta\!\in(\frac{1}{2},\frac{q-1}{q})$ for $q\in\!(2,3]$ is weaker than that of exponent $1/2$. Thus, our convergence results have a wide range of potential applications. Also, the superlinear convergence result of the cost value sequence improves the one got with $p=2$ in [35] for (1), a special case of the optimization problem considered in [35].
@@
(2) Under a local Hölderian error bound of order $\gamma\in\!(\frac{1}{q-1},1]$ on a second-order stationary point set that reduces to the ordinary one used in [23, 47] if $g\equiv 0$, we prove that the iterate sequence converges to a second-order stationary point with a local $Q$-superlinear convergence rate of order $\gamma(q\!-\!1)$, and the convergence of the objective value sequence has the same superlinear rate under an additional lower second-order growth on $F$ around the second-order stationary point. This implies that the iterate and objective value sequences of our inexact CR method converge quadratically to a second-order stationary point and a second-order stationary value, respectively, under a local Lipschitzian error bound. To the best of our knowledge, this is the first practical inexact CR method for (1) to have a quadratic convergence rate without the local strong convexity of $F$, improving the superlinear convergence rate of order $4/3$ obtained in [39] under the KL property of $F$ with exponent $1/2$, and extending the work [47] to nonconvex and nonsmooth composite optimization. It is worth emphasizing that our proof of the local quadratic convergence rate for $q=3$ is not a direct extension of the proofs in [39, 47]. The local error bound condition is much weaker than the strong convexity required in [20], and is basically the same as the one used in [47] when $g\equiv 0$. For the aforementioned quadratic program, though it is not locally strongly convex around any non-minimum stationary points, the local Lipschitzian error bound on the set of stationary points and the KL property of $F$ with exponent $1/2$ are implied by the local upper Lipschitz of $(\partial F)^{-1}$ due to [40, Proposition 1]. We also derive a global complexity bound $O(\epsilon^{-\frac{q}{q-1}})$ in terms of the first-order optimality conditions with the strict continuity of $\nabla^{2}\!f$ on a compact convex set containing the iterate sequence. The complexity bound for $q=3$ matches that of the CR method [37] and its variants [13] for unconstrained smooth optimization, and agrees with the one in [14, Theorem 3.12] if $g=\chi_{C}$ for a closed convex set $C\subset\mathbb{X}$.
@@
(3) We apply the proposed algorithm armed with ZeroFPR [43] as the solver of subproblems to three classes of problem (1), and compare its performance with that of ZeroFPR directly applied to these problems and IRPNM proposed in [30]. Numerical results indicate that our algorithm has the performance matching theoretical results. For the $\ell_{1}$-norm regularized logisitic regression, our CR method returns the optimal values and KKT residuals with better accuracy within $54$ iterations except for data “arcene”. For the nonconvex $\ell_{1}$-norm regularized Student’s t-regression, it yields second-order stationary points with the same objective values and comparable KKT residuals within less iterations; and for the portfolio decision with higher moments, it generates the comparable objective values and better KKT residuals within comparable running time with IRPNM and much less running time than ZeroFPR.
@@ 2 Preliminaries
We first recall the strict continuity and subdifferential continuity of a proper function $h\!:\mathbb{X}\to\overline{\mathbb{R}}$. By [41, Definition 9.1], $h$ is strictly continuous at $\overline{x}\in{\rm dom}\,h$ if there exist a neighborhood $V$ of $\overline{x}$ and a constant $\kappa\geq 0$ such that for any $x^{\prime},x^{\prime\prime}\in V$, $|h(x^{\prime})-h(x^{\prime\prime})|\leq\kappa\|x^{\prime}-x^{\prime\prime}\|$; and it is strictly continuous at $\overline{x}$ relative to ${\rm dom}\,h$ if there exists a neighborhood $V$ of $\overline{x}$ such that $h$ is Lipschitz continuous on ${\rm dom}\,h\cap V$. By [41, Definition 13.28], the function $h$ is subdifferentially continuous at $\overline{x}$ for $\overline{v}$ if $\overline{v}\in\partial h(\overline{x})$ and, whenever $(x^{k},v^{k})\xrightarrow{{\rm gph}\,\partial h}(\overline{x},\overline{v})$, one has $h(x^{k})\to h(\overline{x})$; and when this holds for all $\overline{v}\in\partial h(\overline{x})$, it is said to be subdifferentially continuous at $\overline{x}$.
@@ 2.1 Stationary points of problem (1)
and denote by $\mathcal{X}^{*}$ the set of second-order stationary points. When $g\equiv 0$, this reduces to the one used in [47]. When $g\not\equiv 0$ and $f$ is convex, $\mathcal{X}^{*}=\{x\in\mathcal{S}^{*}\,|\,\nabla^{2}\!f(x)\succeq 0\}=\mathcal{S}^{*}$. When $g\not\equiv 0$ and $f$ is nonconvex, the inclusion $\mathcal{X}^{*}\supset\{x\in\mathcal{S}^{*}\,|\,\nabla^{2}\!f(x)\succeq 0\}$ is generally strict because for a given $x\in\mathcal{S}^{*}$ the set ${\rm prox}_{g}(\mathbb{X})-x$ is much smaller than $\mathbb{X}$; for example, if $g=\chi_{C}$ for a closed convex set $C\subset\mathbb{X}$, ${\rm prox}_{g}(\mathbb{X})-x\subset C-x$.
@@ 2.2 Metric subregularity of residual mapping
(see [27, Definition 3.1]) Let $\mathcal{F}\!:\mathbb{X}\rightrightarrows\mathbb{X}$ be a multifunction. Consider any point $(\overline{x},\overline{y})\in{\rm gph}\,\mathcal{F}$, the graph of $\mathcal{F}$. For a given $\gamma>0$, we say that $\mathcal{F}$ is (metrically) $\gamma$-subregular at $(\overline{x},\overline{y})$ if there exist $\kappa>0$ and $\delta>0$ such that